1(a) a vessel containing an unknown volume of argon gas in cm3 at 300 degree Fahrenheit and 101kpa was inverted and placed in cold ethanol.as the gas contracted,the pressure was quadrupled while the temperature reduced by one -third.what is the final volume of the gas in dm3?
(B) calculate the root mean square of bromine molecule at 23 degree Celsius and 1.00atm.what is the root mean square of the same gas at 23 degree Celsius and 1.5atm
2(a) for the following covalent compounds state the hybridization,bond angle and decide whether their dipole moments will be zero or non-zero (I)IF5(ii) XeF4(iii)XeF2(iv)SF4 present your answer in a tabular form
(b) in one sentence only, account for the differences in bond angle of the following isoelectronic species CH4-109.5degree,NH3-107degree,H2O-105degree
1
Expert's answer
2019-05-01T04:28:49-0400

1(a) a vessel containing an unknown volume of argon gas in cm3 at 300 degree Fahrenheit and 101kpa was inverted and placed in cold ethanol.as the gas contracted,the pressure was quadrupled while the temperature reduced by one -third.what is the final volume of the gas in dm3? 

Solution:

pV=nRT

nR is unchanged; nR=1; T1=422 K; T2=300/3=100 Fahrenheit=310,9 K.

V1=1*422/101000=0,0042 L

V2=1*310,9/404000=0,00077 L

Answer: 0,00077 L

(B) calculate the root mean square of bromine molecule at 23 degree Celsius and 1.00atm.what is the root mean square of the same gas at 23 degree Celsius and 1.5atm 

Urms="\sqrt{\smash[b]{3p/\rho}}"

1."\rho" =p*M/R*T; M(Br2)=2*80=160 grams/mol; R=0,08206 atm*L/mol*K; T=23+273=296;

"\rho" = 1*160/0,08206*296=6,587 grams/L;

Urms = "\sqrt{\smash[b]{3*1/6,587}}" = 0,4554 m/s

2."\rho" =1,5*160/0,08206*296=9,88 grams/L;

Urms="\sqrt{\smash[b]{3*1,5/9,88}}" =0,4555 m/s

Answer: 0,45554 m/s; 0,4555 m/s.

2(b) in one sentence only, account for the differences in bond angle of the following isoelectronic species CH4-109.5degree,NH3-107degree,H2O-105degree

Solution:

The differences in bond angle of the following isoelectronic species CH4-109.5degree,NH3-107degree,H2O-105degree have different amounts of Hydrogen atoms.

(a) for the following covalent compounds state the hybridization,bond angle and decide whether their dipole moments will be zero or non-zero (I)IF5(ii) XeF4(iii)XeF2(iv)SF4 present your answer in a tabular form  

 Solution is the photo.